# Photometry and Spectroscopy of SN 2024pxl: A Luminosity Link Among Type Iax Supernovae

### v1

@@ Abstract
We present extensive ultraviolet to optical photometric and optical to near-infrared (NIR) spectroscopic follow-up observations of the nearby intermediate-luminosity (16.810.19 mag) Type Iax supernova (SN) 2024pxl in NGC 6384. SN 2024pxl exhibits a faster light curve evolution than the high-luminosity members of this class, and slower than low-luminosity events. The observationally well-constrained rise time of 10 days and an estimated synthesized 56Ni mass of 0.03 M⊙, based on analytical modeling of the pseudobolometric light curve, are consistent with models of the weak deflagration of a carbon-oxygen white dwarf. Our optical spectral sequence of SN 2024pxl shows weak Si II lines and spectral evolution similar to other high-luminosity Type Iax SNe, but also prominent early-time C II line, like lower-luminosity Type Iax SNe. The late-time optical spectrum of SN 2024pxl closely matches that of SN 2014dt, and its NIR spectral evolution aligns with those of other well-studied, high-luminosity Type Iax SNe. The spectral-line expansion velocities of SN 2024pxl are at the lower end of the Type Iax SN velocity distribution, and the velocity distribution of iron-group elements compared to intermediate-mass elements suggests that the ejecta are mixed on large scales, as expected in pure deflagration models. SN 2024pxl exhibits characteristics intermediate between those of high-luminosity and low-luminosity Type Iax SNe, further establishing a link across this diverse class.
1 Introduction
Type Iax supernovae (SNe) are the fainter, peculiar cousins of Type Ia SNe, both arising from the explosive thermonuclear fusion of carbon-oxygen (CO) white dwarfs (WDs; for a review, see Jha, 2017). Owing to their homogeneity, “normal” Type Ia SNe serve as standardizable candles in cosmology (Phillips, 1993; Phillips et al., 1999). Type Iax SNe, however, display comparatively lower luminosities, explosion energies, and ejecta velocities (Li et al., 2003; Filippenko, 2003; Foley et al., 2013; Jha, 2017). Their physical properties are also more heterogeneous, with luminosities ranging from mag (Karambelkar et al., 2021) to mag (Stritzinger et al., 2015), and line absorption velocities at maximum light between 2000 and 8000 km s-1 (Foley et al., 2009; Stritzinger et al., 2014). These distinct characteristics suggest that they arise from progenitor systems and/or have explosion mechanisms that differ in detail from normal Type Ia SNe.
@@ 1 Introduction
SN 2024pxl (, , J2000) was first discovered and reported to the Transient Name Server111https://www.wis-tns.org by the BTSbot machine-learning model (Rehemtulla et al., 2024b) on 23 July 2024 at 09:41:18 (UTC dates are used throughout this paper) using data from the Zwicky Transient Facility (ZTF; Bellm et al., 2019a, b; Dekany et al., 2020; Masci et al., 2019; Graham et al., 2019). As part of the BTSbot-nearby program (Rehemtulla et al., 2025), BTSbot immediately and autonomously triggered a target-of-opportunity photometric and spectroscopic request to the spectral energy distribution machine spectrograph (SEDM; Blagorodnova et al., 2018; Kim et al., 2022). SN 2024pxl is located 455 E and 78 N from the nucleus of its host galaxy, NGC 6384. Other designations of this SN are ATLAS24lpk and ZTF24aawrofs. Smartt et al. (2024) classified SN 2024pxl as a Type Iax SN, noting a strong resemblance to the bright Type Iax SN 2005hk one week before maximum light.
@@
This study presents an extensive follow-up campaign of SN 2024pxl, integrating observations from both ground- and space-based telescopes. SN 2024pxl is perhaps the best-sampled Type Iax SN to date in terms of its photometric and spectroscopic data. Kwok et al. (2025) analyze optical and near-infrared (NIR) spectra from this study together with mid-infrared (MIR) data from the James Webb Space Telescope (JWST), providing key insights into the explosion mechanism and progenitor system, complementing the results presented here. Section 2 details the observations and data reduction, and Section 3 describes the distance determination, line-of-sight extinction, and explosion epoch estimation. Section 4 presents the light curve, color evolution, and analytical modeling of the pseudobolometric light curve. In Section 5, we examine the spectral features and line-velocity evolution of SN 2024pxl and compare it with other well-studied Type Iax SNe. Section 6 summarizes the key findings of this study.
2 Observations and Data Reduction
2.1 Photometry
@@
A high-cadence dataset for SN 2024pxl was acquired with both ground- and space-based telescopes as part of an extensive photometric observing campaign initiated on the day of discovery. Photometry in bands was obtained through the Global Supernova Project (GSP) collaboration using the Las Cumbres Observatory (LCO; Brown et al. 2013) 0.4 m and 1 m telescopes. Preprocessing, including bias correction and flat fielding, was handled by the BANZAI pipeline (McCully et al., 2018). Further data reduction was carried out using lcogtsnpipe (Valenti et al., 2016a), a photometric reduction pipeline that uses point-spread-function (PSF) photometry (Stetson, 1987) to calculate zero points, color terms, and extracted magnitudes. Photometry in bands is reported in Vega magnitudes (Stetson, 2000), while band data are presented in AB magnitudes (Oke & Gunn, 1983), calibrated against Sloan Digital Sky Survey (SDSS) sources (Smith et al., 2002). Given the proximity of SN 2024pxl to its host galaxy, host contamination was mitigated by subtracting template images acquired on March 7, 2018 — before the explosion. These templates, originally taken for SN 2017drh (which exploded in the same galaxy), were subtracted using the PyZOGY algorithm (Zackay et al., 2016; Guevel & Hosseinzadeh, 2017), integrated within the lcogtsnpipe pipeline.
@@
Additional imaging of SN 2024pxl was obtained in the bands with the 1 m Nickel telescope at Lick Observatory. The images were calibrated using bias and sky flat-field frames following standard procedures. PSF photometry was performed and calibrated relative to Pan-STARRS photometric standards (Flewelling et al., 2020). Optical photometry of SN 2024pxl is presented in LABEL:tab:photometric_observational_log_2024pxl.
@@
SN 2024pxl was also observed in the z band with the RC80 and BRC80 robotic telescopes at Piszkesteto station of Konkoly Observatory and at Baja Observatory of University of Szeged, Hungary (a complete BgVriz light curve will be presented by Banhidi et al., in prep.). The magnitudes were computed via aperture photometry and tied to Pan-STARRS photometry of local stars, including a color term in the transformation. The z-band photometry of SN 2024pxl is provided in LABEL:tab:z_band_photometry_24pxl.
@@
The Neil Gehrels Swift Observatory (Swift Gehrels et al., 2004) observed SN 2024pxl with its Ultra-Violet/Optical Telescope (UVOT; Roming et al. 2005) in both ultraviolet (UV) and optical filters. We reduced the UVOT images using the High-Energy Astrophysics Software (HEASoft222https://heasarc.gsfc.nasa.gov/docs/software/heasoft/). A circular source region centered at the position of the SN with a radius of 3 was used for aperture photometry. We measured the background contribution from a circular region (aperture radius of 5) that is not contaminated by any other sources. Zeropoints for photometry were chosen from Breeveld et al. (2010) with time-dependent sensitivity corrections updated in 2020. Table 5 details the photometric observations of SN 2024pxl collected by Swift.
2.2 Spectroscopy
2.2.1 Optical Spectra
@@
We have compiled an extensive spectral dataset for SN 2024pxl using multiple ground-based telescopes in the optical and near-infrared (NIR) domains. Spectra of SN 2024pxl obtained with the Double Beam Spectrograph (DBSP; Oke & Gunn 1982) mounted on the 5 m Hale Telescope at Palomar Observatory were reduced using DBSP-DRP (Mandigo-Stoba et al., 2021, 2022a, 2022b) and PypeIt (Prochaska et al., 2020). Spectra acquired with the Wide Field Spectrograph (WiFeS; Dopita et al. 2007, 2010) on Australian National University (ANU) 2.3 m telescope at Siding Spring Observatory were reduced with the data-reduction pipeline PyWiFeS (Childress et al., 2014; Carr et al., 2024). Spectroscopic observations were also triggered with the FLOYDS spectrograph on the 2 m Faulkes Telescope North and South (FTN and FTS; Brown et al. 2013) through the GSP collaboration. The spectra were reduced with the floydsspec333https://www.authorea.com/users/598/articles/6566 pipeline, using standard reduction techniques.
@@
Gemini IRAF444IRAF is distributed by the National Optical Astronomy Observatories, operated by the Association of Universities for Research in Astronomy, Inc., under a cooperative agreement with the National Science Foundation. packages were used to reduce the spectra of SN 2024pxl from the Gemini Multi-Object Spectrograph mounted at Gemini North (GMOS-N; Hook et al. 2004) and Gemini South (GMOS-S) Observatories. Two spectra of SN 2024pxl were acquired using the SPectrograph for the Rapid Acquisition of Transients (SPRAT; Piascik et al. 2014), mounted on the Liverpool Telescope (LT; Steele & et al. 2004). Reduction of the SPRAT spectra was carried out using the SPRAT pipeline.
@@
SN 2024pxl was also observed with the Robert Stobie Spectrograph (RSS) attached to the 9.2 m Southern African Large Telescope (SALT). All RSS spectra were reduced using a custom pipeline based on standard Pyraf (Science Software Branch at STScI, 2012) routines and the PySALT package (Crawford et al., 2010). A few spectra of SN 2024pxl were taken from the Marcario Low-Resolution Spectrograph (LRS) mounted on Hobby–Eberly Telescope (HET) located in McDonald Observatory (Ramsey et al., 1998). These were reduced following standard IRAF packages.
@@
As part of our observing campaign, we also utilized the Hanle Faint Object Spectrograph and Camera (HFOSC) mounted on the Himalayan Chandra Telescope (HCT; Prabhu & Anupama 2010). The HCT spectra were reduced using standard IRAF routines.
@@
SN 2024pxl was also observed with Binospec on the MMT telescope (Fabricant et al., 2019); these spectra were reduced using the Binospec IDL pipeline (Kansky et al., 2019). A spectrum was taken with the Low Resolution Imaging Spectrometer (LRIS; Oke et al., 1995) on the Keck I 10 m telescope, which was reduced in a standard way using the LPipe pipeline (Perley, 2019). Many spectra of SN 2024pxl were obtained with the Kast double spectrograph (Miller & Stone, 1993) on the Shane 3 m telescope at Lick Observatory, either at low airmass or with the slit aligned along the parallactic angle to minimize the effects of atmospheric dispersion (Filippenko, 1982). The reduction of all Kast spectra was performed following standard procedures555https://github.com/msiebert1/UCSC_spectral_pipeline,666https://github.com/ishivvers/TheKastShiv as outlined by Siebert et al. (2019) and Silverman et al. (2012).
@@
In our observing campaign, we also acquired data from the Gran Telescopio Canarias (GTC) with OSIRIS and EMIR, and reduced the spectra using a dedicated pipeline based on PypeIt (Prochaska et al., 2020a, b, 2021; Galbany et al., 2025). PypeIt was also used to reduce the spectra of SN 2024pxl gathered with the 2.56 m Nordic Optical Telescope (NOT) deploying the Alhambra Faint Object Spectrograph and Camera (ALFOSC). We collected one spectrum of SN 2024pxl with the XShooter spectrograph mounted on the Very Large Telescope (Vernet et al., 2011); the EsoReflLex pipeline (Freudling et al., 2013), including the XShooter module, was used to reduce it.

### v2

@@ Abstract
We present extensive ultraviolet to optical photometric and optical to near-infrared (NIR) spectroscopic follow-up observations of the nearby intermediate-luminosity ( mag) Type Iax supernova (SN) 2024pxl in NGC 6384. SN 2024pxl exhibits a faster light curve than the high-luminosity members of this class, and slower than low-luminosity events. The observationally well-constrained rise time of 11 days and an estimated synthesized 56Ni mass of 0.03 M⊙, based on analytical modeling of the integrated spectral energy distribution light curve, are consistent with models of the weak deflagration of a carbon-oxygen white dwarf. Our optical spectral sequence of SN 2024pxl shows weak Si II lines and spectral evolution similar to other high-luminosity Type Iax SNe, but also a prominent early-time C II line, like lower-luminosity Type Iax SNe. The late-time optical spectrum of SN 2024pxl closely matches that of SN 2014dt, and its NIR spectral evolution aligns with that of other well-studied, high-luminosity Type Iax SNe. The spectral-line expansion velocities of SN 2024pxl are at the lower end of the Type Iax SN velocity distribution, and the velocity distribution of iron-group elements compared to intermediate-mass elements suggests that the ejecta are mixed on large scales, as expected in pure deflagration models. SN 2024pxl exhibits characteristics intermediate between those of high-luminosity and low-luminosity Type Iax SNe, further establishing a link across this diverse class.
I Introduction
Type Iax supernovae (SNe) are the fainter, peculiar cousins of Type Ia SNe, both arising from the explosive thermonuclear fusion of carbon-oxygen (CO) white dwarfs (WDs; for a review, see Jha, 2017). Owing to their homogeneity, “normal” Type Ia SNe serve as standardizable candles in cosmology (Phillips, 1993; Phillips et al., 1999). Type Iax SNe, however, display comparatively lower luminosities, explosion energies, and ejecta velocities (Li et al., 2003; Filippenko, 2003; Foley et al., 2013; Jha, 2017). Their physical properties are also more heterogeneous, with luminosities ranging from mag (Karambelkar et al., 2021) to mag (Stritzinger et al., 2015), and line absorption velocities at maximum light between 2000 and 8000 km s-1 (Foley et al., 2009; Stritzinger et al., 2014). These distinct characteristics suggest that they arise from different progenitor systems and also have explosion mechanisms that differ in detail from those of normal Type Ia SNe.
@@ I Introduction
SN 2024pxl (, , J2000) was first discovered and reported to the Transient Name Server111https://www.wis-tns.org by the BTSbot machine-learning model (Rehemtulla et al., 2024b) on 23 July 2024 at 09:41:18 (UTC dates are used throughout this paper) using data from the Zwicky Transient Facility (ZTF; Bellm et al., 2019a, b; Dekany et al., 2020; Masci et al., 2019; Graham et al., 2019). As part of the BTSbot-nearby program (Rehemtulla et al., 2025), BTSbot triggered a target-of-opportunity photometric and spectroscopic request to the spectral energy distribution machine spectrograph (SEDM; Blagorodnova et al., 2018; Kim et al., 2022). SN 2024pxl is located 455 E and 78 N from the nucleus of its host galaxy, NGC 6384. Other designations of this SN are ATLAS24lpk and ZTF24aawrofs. Smartt et al. (2024) classified SN 2024pxl as a Type Iax SN, noting a strong resemblance to the bright Type Iax SN 2005hk one week before maximum light.
@@
This study presents an extensive follow-up campaign of SN 2024pxl, integrating observations from both ground- and space-based telescopes. SN 2024pxl is perhaps the best-sampled Type Iax SN to date in terms of its photometric and spectroscopic data. Kwok et al. (2025) analyze optical and near-infrared (NIR) spectra from this study together with mid-infrared (MIR) data from the James Webb Space Telescope (JWST), providing important clues about the explosion mechanism and progenitor system, complementing the results presented here. Section II details the observations and data reduction, and Section III describes the distance determination, line-of-sight extinction, and explosion epoch estimation. Section IV presents the light curve, color evolution, and analytical modeling of the integrated spectral energy distribution (SED) light curve. In Section V, we examine the spectral features and line-velocity evolution of SN 2024pxl and compare it with other well-studied Type Iax SNe. Section VI summarizes the key findings of this study.
II Observations and Data Reduction
II.1 Photometry
@@
A high-cadence dataset for SN 2024pxl was acquired with both ground- and space-based telescopes as part of an extensive photometric observing campaign initiated on the day of discovery. Photometry in bands was obtained through the Global Supernova Project (GSP) collaboration using the Las Cumbres Observatory (LCO; Brown et al. 2013) 0.4 m and 1 m telescopes. Preprocessing, including bias correction and flat fielding, was handled by the BANZAI pipeline (McCully et al., 2018). Further data reduction was carried out using lcogtsnpipe (Valenti et al., 2016a), a photometric reduction pipeline that uses point-spread-function (PSF) photometry (Stetson, 1987a) to calculate zero points, color terms, and extracted magnitudes. Photometry in bands is reported in Vega magnitudes (Landolt, 1992a), while band data are presented in AB magnitudes (Oke & Gunn, 1983), calibrated against Sloan Digital Sky Survey (SDSS) sources (Smith et al., 2002). Given the proximity of SN 2024pxl to its host galaxy, host contamination was mitigated by subtracting template images acquired on March 7, 2018 — before the explosion. These templates, originally taken for SN 2017drh (which exploded in the same galaxy), were subtracted using the PyZOGY algorithm (Zackay et al., 2016; Guevel & Hosseinzadeh, 2017), integrated within the lcogtsnpipe pipeline.
@@
Additional imaging of SN 2024pxl was obtained in bands with the 1 m Nickel telescope at Lick Observatory. The images were calibrated using bias and sky flat-field frames following standard procedures. PSF photometry was performed and calibrated relative to Pan-STARRS1 photometric standards (Flewelling et al., 2020).
Early-time photometric observations of SN 2024pxl were also obtained through the Distance Less Than 40 Mpc (DLT40) survey (Tartaglia et al., 2018), using the PROMPT-MO 0.4 m telescope operated via the Skynet Robotic Telescope Network (Reichart et al., 2005). These observations were carried out in a broad, unfiltered “Open” mode, and the resulting data were transformed to the SDSS band following the calibration and reduction procedures described by Tartaglia et al. (2018).
Multiband photometry was further acquired using the 0.76 m Katzman Automatic Imaging Telescope (KAIT; Filippenko et al., 2001) at Lick Observatory in the BVRI bands. All images were reduced with a custom pipeline222https://github.com/benstahl92/LOSSPhotPypeline described by Ganeshalingam et al. (2010) and Stahl et al. (2019a). Host-galaxy contamination was removed via image subtraction, using a pre-explosion template obtained on 2018 June 22. PSF photometry was performed with DAOPHOT (Stetson, 1987b) from the IDL Astronomy User’s Library333http://idlastro.gsfc.nasa.gov/. Photometric calibration was achieved using local standard stars from the Pan-STARRS1 Surveys (Schlafly et al., 2012), with magnitudes transformed into the Landolt standard system (Landolt, 1992b) using the relations from Tonry et al. (2012).
In addition, observations of SN 2024pxl were conducted in the bands using the Direct 4k 4k camera on the 1 m Henrietta Swope telescope at Las Campanas Observatory, Chile. A full description of the data-reduction procedure is provided by Kilpatrick et al. (2018).
SN 2024pxl was observed with the Dark Energy Camera (DECam) in the bands as part of the Young Supernova Experiment (YSE DECam; NOIRLab propID 2023A-237157). Data reduction, image differencing, and forced photometry were performed using photpipe (Rest et al., 2014). Templates included public DECam Local Volume Exploration Survey (DELVE; Drlica-Wagner et al. 2021) pre-explosion images from 2023 September 9 (, ) and 2023 April 7 (), as well as a public BLINK (Billion Lines INdexing in a clicK; Kamennoff et al. 2012) image from 2018 April 13 ().
@@
SN 2024pxl was also observed in the z band with the RC80 and BRC80 robotic telescopes at Piszkesteto station of Konkoly Observatory and at Baja Observatory of University of Szeged, Hungary (a complete BgVriz light curve will be presented by Banhidi et al., in prep.). The magnitudes were computed via aperture photometry and tied to Pan-STARRS photometry of local stars, including a color term in the transformation.
The multiband photometry from the different instruments is consistent within the quoted uncertainties, with no significant systematic offsets observed. Optical photometry of SN 2024pxl is presented in Table LABEL:tab:photometric_observational_log_2024pxl. The z-band photometry of SN 2024pxl is provided in LABEL:tab:z_band_photometry_24pxl.
@@
The Neil Gehrels Swift Observatory (Swift Gehrels et al., 2004) observed SN 2024pxl with its Ultra-Violet/Optical Telescope (UVOT; Roming et al. 2005) in both ultraviolet (UV) and optical filters. We reduced the UVOT images using the High-Energy Astrophysics Software (HEASoft444https://heasarc.gsfc.nasa.gov/docs/software/heasoft/). A circular source region centered at the position of the SN with a radius of 3 was used for aperture photometry. We measured the background contribution from a circular region (aperture radius of 5) that is not contaminated by any other sources. Zeropoints for photometry were chosen from Breeveld et al. (2010) with time-dependent sensitivity corrections updated in 2020. Table 5 details the photometric observations of SN 2024pxl collected by Swift.
II.2 Spectroscopy
II.2.1 Optical Spectra
@@
We have compiled an extensive spectral dataset for SN 2024pxl using multiple ground-based telescopes in the optical and near-infrared (NIR) domains. Spectra of SN 2024pxl obtained with the Double Beam Spectrograph (DBSP; Oke & Gunn 1982) mounted on the 5 m Hale Telescope at Palomar Observatory were reduced using DBSP-DRP (Mandigo-Stoba et al., 2021, 2022a, 2022b) and PypeIt (Prochaska et al., 2020). Spectra acquired with the Wide Field Spectrograph (WiFeS; Dopita et al. 2007, 2010) on Australian National University (ANU) 2.3 m telescope at Siding Spring Observatory were reduced with the data-reduction pipeline PyWiFeS (Childress et al., 2014; Carr et al., 2024). Spectroscopic observations were also triggered with the FLOYDS spectrograph on the 2 m Faulkes Telescope North and South (FTN and FTS; Brown et al. 2013) through the GSP collaboration. The spectra were reduced with the floydsspec555https://www.authorea.com/users/598/articles/6566 pipeline, using standard reduction techniques.
@@
Gemini IRAF666IRAF is distributed by the National Optical Astronomy Observatories, operated by the Association of Universities for Research in Astronomy, Inc., under a cooperative agreement with the National Science Foundation. packages were used to reduce the spectra of SN 2024pxl from the Gemini Multi-Object Spectrograph mounted at Gemini North (GMOS-N; Hook et al. 2004) and Gemini South (GMOS-S) Observatories. Two spectra of SN 2024pxl were acquired using the SPectrograph for the Rapid Acquisition of Transients (SPRAT; Piascik et al. 2014), mounted on the Liverpool Telescope (LT; Steele & et al. 2004). Reduction of the SPRAT spectra was carried out using the SPRAT pipeline.
@@
SN 2024pxl was observed with the Robert Stobie Spectrograph (RSS) attached to the 9.2 m Southern African Large Telescope (SALT). All RSS spectra were reduced using a custom pipeline based on standard Pyraf (Science Software Branch at STScI, 2012) routines and the PySALT package (Crawford et al., 2010).
A few spectra of SN 2024pxl were taken with the Low Resolution Spectrograph 2 (LRS2) (Chonis et al., 2014) mounted on the Hobby-Eberly Telescope (HET) (Ramsey et al., 1998) located at McDonald Observatory. LRS2 is a two-beam spectrograph covering the blue (LRS2-B) and the red (LRS2-R) part of the optical spectrum with an average resolution of . The data were reduced by the Panacea pipeline777https://github.com/grzeimann/Panacea.
@@
As part of our observing campaign, we have utilized the Hanle Faint Object Spectrograph and Camera (HFOSC) mounted on the Himalayan Chandra Telescope (HCT; Prabhu & Anupama 2010). The HCT spectra were reduced using standard IRAF routines.
@@
SN 2024pxl was observed with Binospec on the MMT telescope (Fabricant et al., 2019); these spectra were reduced using the Binospec IDL pipeline (Kansky et al., 2019). A spectrum was taken with the Low Resolution Imaging Spectrometer (LRIS; Oke et al., 1995) on the Keck I 10 m telescope, which was reduced in a standard way using the LPipe pipeline (Perley, 2019). Many spectra of SN 2024pxl were obtained with the Kast double spectrograph (Miller & Stone, 1993) on the Shane 3 m telescope at Lick Observatory, either at low airmass or with the slit aligned along the parallactic angle to minimize the effects of atmospheric dispersion (Filippenko, 1982). The reduction of all Kast spectra was performed following standard procedures888https://github.com/msiebert1/UCSC_spectral_pipeline,999https://github.com/ishivvers/TheKastShiv as outlined by Siebert et al. (2019) and Silverman et al. (2012).
@@
In our observing campaign, we acquired data from the Gran Telescopio Canarias (GTC) with OSIRIS and EMIR, and reduced the spectra using a dedicated pipeline based on PypeIt (Prochaska et al., 2020a, b, 2021; Galbany et al., 2025). PypeIt was also used to reduce the spectra of SN 2024pxl gathered with the 2.56 m Nordic Optical Telescope (NOT) deploying the Alhambra Faint Object Spectrograph and Camera (ALFOSC). We collected one spectrum of SN 2024pxl with the XShooter spectrograph mounted on the Very Large Telescope (Vernet et al., 2011); the EsoReflLex pipeline (Freudling et al., 2013), including the XShooter module, was used to reduce it.